### v1

# Exploiting Inter-Sample Correlation and Intra-Sample Redundancy for Partially Relevant Video Retrieval

Abstract.
Partially Relevant Video Retrieval (PRVR) aims to retrieve the target video that is partially relevant to the text query. The primary challenge in PRVR arises from the semantic asymmetry between textual and visual modalities, as videos often contain substantial content irrelevant to the query. Existing methods coarsely align paired videos and text queries to construct the semantic space, neglecting the critical cross-modal dual nature inherent in this task: inter-sample correlation and intra-sample redundancy. To this end, we propose a novel PRVR framework to systematically exploit these two characteristics. Our framework consists of three core modules. First, the Inter Correlation Enhancement (ICE) module captures inter-sample correlation by identifying semantically similar yet unpaired text queries and video moments, combining them to form pseudo-positive pairs for more robust semantic space construction. Second, the Intra Redundancy Mining (IRM) module mitigates intra-sample redundancy by mining redundant video moment features and treating them as hard negative samples, thereby encouraging the model to learn more discriminative representations. Finally, to reinforce these modules, we introduce the Temporal Coherence Prediction (TCP) module, which enhances feature discrimination by training the model to predict the original temporal order of randomly shuffled video frames and moments. Extensive experiments on three datasets demonstrate the superiority of our approach compared to previous methods, achieving state-of-the-art results.
Cross-Modal Retrieval, Video Understanding, Multimodal Learning, Multimedia Applications
copyright: acmlicensedjournalyear: 2017doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; 2017; ccs: Information systems Multimedia and multimodal retrievalccs: Information systems Video search
1. Introduction
Figure 1. In Partially Relevant Video Retrieval (PRVR), video-text samples exhibit the inherent cross-modal dual nature: (a) Inter-sample correlation: The video contains certain moments that are semantically correlated to other unpaired text queries. (b) Intra-sample redundancy: Apart from the target moment, other redundant moments in the video are irrelevant to the paired text query.
With the explosion of online video content, users increasingly seek efficient ways to find specific videos within vast video collections. While conventional video retrieval methods focus on identifying videos that fully match a given query, many real-world scenarios require identifying videos that contain only partial moments relevant to a user’s descriptive text query. This task is formally known as Partially Relevant Video Retrieval (PRVR).
The fundamental challenge in PRVR stems from the inherent semantics asymmetry between text and video modalities. While a video comprises multiple moments, only a subset may exhibit relevance to a given text query. Existing methods (Dong et al., 2022a; Jiang et al., 2023; Dong et al., 2023; Wang et al., 2024b; Yin et al., 2024) approach this task from a limited perspective, primarily adopting conventional alignment constraints such as triplet ranking (Faghri et al., 2018) and InfoNCE (Oord et al., 2018) to establish coarse associations between annotated paired text query and the target videos. These approaches may be effective when the semantic content of the two modalities is largely equivalent, as in the traditional text-to-video retrieval tasks.
However, these methods overlook the unique cross-modal dual nature in the PRVR task: namely inter-sample correlation and intra-sample redundancy. As illustrated in Figure 1, the video modality inherently shows richer semantics than the corresponding text query. Moments not explicitly related to the query may still contain valuable visual semantics that correlate with other unpaired queries in the dataset. Conversely, the video modality also suffers from semantic redundancy, as only a limited subset of its content aligns with the given textual descriptions (Dong et al., 2022a). This results in a substantial portion of video content being irrelevant to the text query, which may not only fail to contribute to retrieval accuracy but could actively hinder the proper alignment between text queries and their corresponding video moments.
Based on these observations, our approach aims to enhance PRVR by simultaneously leveraging inter-sample correlations and mitigating intra-sample redundancy. To harness inter-sample correlations, we propose to fully utilize the rich semantics presented in videos. Technically, our method identifies and exploits underlying semantic relationships between unpaired video moments and text queries within the training set. This approach stems from the key insight that while certain video moments may not directly correspond to their paired text descriptions, they often share meaningful semantic connections with other queries in the dataset. By establishing these additional cross-modal associations, we enhance the construction of a more robust and discriminative semantic space.
To address intra-sample redundancy, our framework explicitly mitigates the interference caused by irrelevant video moments in video-text matching. We propose to treat redundant moment features as hard negative samples during model training. These redundant features, while semantically distinct from the target moment, exhibit high similarity as they originate from the same video. By forcing the model to distinguish these challenging negative samples and true positive matches, we encourage the learning of more precise and discriminative visual representations that better align with their corresponding textual descriptions.
In this paper, we introduce a novel framework for PRVR that systematically exploits the cross-modal dual nature of video-text relationships through two key characteristics: inter-sample correlation and intra-sample redundancy. The proposed framework comprises three key components: (1) The Inter Correlation Enhancement module is designed to fully exploit the inter-sample correlation of video-text modalities. It computes the cross-modal similarity between video moments and unpaired text queries. High-similarity pairs are identified and utilized as pseudo-positive samples, which are then incorporated into the alignment objective to construct a more comprehensive and discriminative cross-modal embedding space. (2) To address intra-sample redundancy, the Intra Redundancy Mining module extracts semantically redundant video moments through joint analysis of text features, global video features, and local moment features. These redundant features serve as hard negative samples during training, forcing the model to strengthen alignment between text and relevant video content and improve discrimination against irrelevant visual semantics within the same video. (3) Complementing the above modules, the Temporal Coherence Prediction module enhances temporal feature discrimination through an auxiliary self-supervised task. The model learns to predict the original temporal order of randomly shuffled video frames/moments, thereby developing a more robust understanding of video semantics and temporal structure.
Extensive experiments on three large-scale datasets, TVR (Lei et al., 2020), ActivityNet Captions (Krishna et al., 2017), and Charades-STA (Gao et al., 2017), verify the superior performance and robustness of our proposed method. In summary, our principal contributions are summarized as follows:
  • We introduce the Inter Correlation Enhancement module, which identifies high-similarity pairs between video moments and unpaired text queries. These pseudo-positive pairs are incorporated to enrich the cross-modal semantic space construction.
  • We propose the Intra Redundancy Mining module to extract and leverage redundant video moment features as hard negative samples. It forces the model to develop more precise text-video alignment by distinguishing relevant semantics from redundant visual content within the same video.
  • In the Temporal Coherence Prediction module, we design a self-supervised auxiliary task. It enhances temporal feature discrimination by requiring the model to predict the original temporal order of shuffled video frames/moments. This results in more robust and discriminative video features.
2. Related Work
2.1. Partially Relevant Video Retrieval
Partially Relevant Video Retrieval (PRVR) is to retrieve untrimmed videos that are only partially semantically related to the text query. Unlike previous video retrieval tasks, videos retrieved by PRVR do not necessarily exhibit complete semantic alignment with the textual query, often containing substantial redundant information. Specifically, MS-SL (Dong et al., 2022a) is the first to propose the PRVR task. MS-SL learns video features at both the clip-scale and frame-scale. PEAN (Jiang et al., 2023) introduces a Gaussian-based pooling strategy to improve the model’s holistic understanding of events. DL-DKD (Dong et al., 2023) distills CLIP’s (Radford et al., 2021) prior knowledge of image-text matching and mitigates the domain gap. GMMFormer (Wang et al., 2024b) increases information density in video encoding by adopting Gaussian windows in Transformer blocks (Vaswani et al., 2017). BGM-Net (Yin et al., 2024) proposes a new loss function to facilitate multimodal alignment based on unimodal similarities. However, these works neglect the unique cross-modal dual nature in PRVR, leading to limited semantic space construction. In contrast, we aim to exploit the inter-sample correlation and intra-sample redundancy, resulting in improved retrieval accuracy.
Figure 2. Overview of the proposed framework. We systematically leverage the cross-modal dual nature in PRVR, namely inter-sample correlation and intra-sample redundancy, to construct a more discriminative cross-modal semantic space. The framework comprises three key components: (a) Inter Correlation Enhancement Module: This component analyzes cross-modal correlations by identifying high-similarity pairs between video moments and unpaired text queries. These pseudo-positive pairs are incorporated during training to enrich the semantic space construction. (b) Intra Redundancy Mining Module: The module extracts and utilizes redundant video moments as hard negative samples. By learning to distinguish these hard negatives, the model develops enhanced capability to focus on query-relevant visual semantics. (c) Temporal Coherence Prediction Module: Designed to complement the other components, this module improves visual feature discrimination through a self-supervised sequence prediction task, where the model predicts the original temporal order of shuffled video frames/moments.
2.2. Video Moment Retrieval
Video Moment Retrieval (VMR) aims to extract the start/end time boundaries of semantically relevant moments from an untrimmed video with text queries. Existing methods can be generally classified into proposal-based (Yuan et al., 2019; Zhang et al., 2020a, 2021a; Wang et al., 2022; Liu et al., 2023; Fang et al., 2024) and proposal-free (Lu et al., 2019; Zhang et al., 2020b; Zhou et al., 2021; Hao et al., 2022; Zheng et al., 2023; Dong et al., 2024) approaches. Proposal-based methods first generate a set of candidate moment proposals and subsequently identify the best-matching proposal. In contrast, proposal-free methods directly predict the start and end timestamps by integrating video and query features, thus eliminating the need for explicit proposals. Departing from these conventional methods, recent studies (Lei et al., 2021; Cao et al., 2021; Liu et al., 2022; Moon et al., 2023; Lin et al., 2023; Jung et al., 2025) also utilize the architecture of DETR (Carion et al., 2020) to address VMR as a set prediction task. Unlike VMR, PRVR focuses on retrieving untrimmed videos instead of temporal locations of moments, thus the temporal annotations are not provided in PRVR.
2.3. Text-to-Video Retrieval
In recent years, Text-to-Video Retrieval (T2VR) has attracted increasing attention. A common approach (Chen et al., 2020; Li et al., 2019; Dong et al., 2019, 2021, 2022b) involves retrieving pre-trimmed videos by comparing cross-modal similarities between textual queries and video content. To achieve this, many studies (Wu et al., 2023; Li et al., 2023; Wang et al., 2024a) have proposed methods that project features from different modalities into a shared embedding space. However, these approaches often neglect the fact that real-world videos may not consistently correspond to a single topic, and typically certain all moments within a video are relevant to the textual query.
3. Proposed Method
In this section, we elaborate on the technical details of our proposed framework, which is illustrated in Figure 2. We first delineate the problem formulation in Section 3.1. Then the base model architecture is expounded in Section 3.2. The three key modules, namely Inter Correlation Enhancement module (ICE), Intra Redundancy Mining module (IRM), and Temporal Coherence Prediction module (TCP) are presented in Section 3.3, Section 3.4, and Section 3.5, respectively. Finally, the training objective and model inference are detailed in Section 3.6.
3.1. Problem Formulation
Partially relevant video retrieval (PRVR) aims to retrieve videos containing a moment semantically relevant to a textual query, from a large corpus of untrimmed videos. Unlike conventional text-to-video retrieval, the videos are untrimmed and much longer, and text queries typically correspond to a small portion of a video. Note that the start/end timestamps of moments are unavailable in PRVR.
3.2. Base Model
3.2.1. Text Representation
Given a sentence containing $N$ words, following previous works (Dong et al., 2022a; Wang et al., 2024b; Yin et al., 2024), we first employ the pre-trained RoBERTa (Liu et al., 2019) to extract word embeddings. Then we utilize a fully-connected (FC) layer to project the word embeddings into a lower-dimensional space. We further adopt a standard Transformer layer (Vaswani et al., 2017) as the text encoder. The encoded word-level features are denoted as $Q=\left\{q_{i}\right\}_{i=1}^{N}\in\mathbb{R}^{N\times D}$, where $D$ is the feature dimension. Finally, we apply the additive attention mechanism (Bahdanau et al., 2015) on $Q$ to obtain the aggregated sentence-level feature $q\in\mathbb{R}^{D}$:
(1) $\alpha^{q}={Softmax}\left(w^{q}\cdot Q^{\top}\right)\in\mathbb{R}^{1\times N},~{}q=\sum_{i=1}^{N}\alpha_{i}^{q}\times q_{i},$
where $w^{q}\in\mathbb{R}^{1\times D}$ is a learnable weight.
3.2.2. Video Representation
Given an untrimmed video, we first adopt a pre-trained 2D or 3D CNN (Carreira and Zisserman, 2017) as the visual backbone to extract $T_{f}$ frame features. Following previous methods (Dong et al., 2022a; Wang et al., 2024b; Yin et al., 2024; Jiang et al., 2023), we employ a coarse-to-fine manner to capture the temporal information of videos by two distinct branches. Technically, the moment-level branch measures the local moment-text similarity while the video-level branch models the global video-text similarity.
For the moment-level branch, we first condense the $T_{f}$ frame features into $T_{m}$ moment features by mean pooling over consecutive frame features. We then use an FC layer to reduce dimension. For the moment encoder, a standard Transformer encoder (Vaswani et al., 2017) is adopted to get contextual moment features $V_{m}=\left\{m_{i}\right\}_{i=1}^{T_{m}}\in\mathbb{R}^{T_{m}\times D}$.
For the video-level branch, the frame features are fed into an FC layer, followed by the video encoder, which is a standard Transformer encoder. The encoded frame features are denoted as $V_{f}=\left\{f_{i}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$. Next, we employ the the additive attention mechanism (Bahdanau et al., 2015) on $V_{f}$ to obtain the aggregated video-level feature $v\in\mathbb{R}^{D}$:
(2) $\alpha^{v}=\operatorname{\textit{Softmax}}\left(w^{v}\cdot V_{f}^{\top}\right)\in\mathbb{R}^{1\times T_{f}},~{}v=\sum_{i=1}^{T_{f}}\alpha_{i}^{v}\times f_{i},$
where $w^{v}\in\mathbb{R}^{1\times D}$ is a learnable weight.
3.2.3. Similarity Learning
Given a video-text pair $\mathcal{V}$ and $\mathcal{T}$, we first compute the similarity between the text representation and video representations from the two aforementioned branches. Concretely, the video-level similarity between the video feature $v$ and the sentence feature $q$ is measured by the cosine similarity:
(3) $\mathcal{S}_{v}\left(\mathcal{V},\mathcal{T}\right)=\operatorname{\textit{cos}}\left(v,q\right)=\frac{v^{\top}q}{\left\|v\right\|\cdot\left\|q\right\|}.$
Next, we compute the moment-level similarity between moment features $V_{m}$ and the sentence feature $q$ through the max-pooling operation and identify the target moment feature $m^{t}\in\mathbb{R}^{D}$:
(4) $\mathcal{S}_{m}\left(\mathcal{V},\mathcal{T}\right)=\operatorname{\textit{max}}\{\operatorname{\textit{cos}}\left(m_{1},q\right),\cdots,\operatorname{\textit{cos}}\left(m_{T_{m}},q\right)\}=\operatorname{\textit{cos}}\left(m^{t},q\right).$
For training, we adopt the triplet ranking (Faghri et al., 2018) and InfoNCE (Oord et al., 2018) losses, which is a common approach following (Dong et al., 2022a; Jiang et al., 2023; Dong et al., 2023; Wang et al., 2024b; Yin et al., 2024):
(5) $\mathcal{L}^{{trip}=}\frac{1}{n}\sum_{(\mathcal{V},\mathcal{T})\in\mathcal{B}}[$ $max\left(0,m+\mathcal{S}\left(\mathcal{V},\mathcal{T}^{-}\right)-\mathcal{S}(\mathcal{V},\mathcal{T})\right)$
$+$ $max(0,m+\mathcal{S}\left(\mathcal{V}^{-},\mathcal{T}\right)-\mathcal{S}(\mathcal{V},\mathcal{T}))],$
$\mathcal{L}^{nce}=-\frac{1}{n}\sum_{(\mathcal{V},\mathcal{T})\in\mathcal{B}}{\left[\log\left(\frac{\mathcal{S}(\mathcal{V},\mathcal{T})}{\sum_{\mathcal{T}_{i}}\mathcal{S}\left(\mathcal{V},\mathcal{T}_{i}\right)}\right)\right.}\left.+\log\left(\frac{\mathcal{S}(\mathcal{V},\mathcal{T})}{\sum_{\mathcal{V}_{i}}\mathcal{S}\left(\mathcal{V}_{i},\mathcal{T}\right)}\right)\right],$
where $m$ is the margin constant, $n$ is the size of the mini-batch $\mathcal{B}$, $\mathcal{T}^{-}$ and $\mathcal{V}^{-}$ are negative samples from $\mathcal{B}$, $\mathcal{T}_{i}$ and $\mathcal{V}_{i}$ denote all text and video samples. The training loss of the base model is:
(7) $\mathcal{L}_{base}=\mathcal{L}^{trip}_{v}+\mathcal{L}^{trip}_{m}+\lambda_{1}\mathcal{L}^{nce}_{v}+\lambda_{2}\mathcal{L}^{nce}_{m},$
where $\mathcal{L}^{trip}_{v}$ and $\mathcal{L}^{trip}_{m}$ denote triplet losses using the video-level similarity $\mathcal{S}_{v}$ and moment-level similarity $\mathcal{S}_{m}$ as $\mathcal{S}$, and accordingly for $\mathcal{L}^{nce}_{v}$ and $\mathcal{L}^{nce}_{m}$. $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters to balance losses.
3.3. Inter Correlation Enhancement
The PRVR task presents a fundamental semantic asymmetry, where textual queries only partially describe the content of target videos, leaving the visual modality inherently richer in semantic information. Our key observation reveals that while certain video moments may not align with their originally paired text annotations, they may exhibit strong semantic correspondence with other unpaired textual descriptions in the dataset as shown in Figure 1. This inter-sample correlation represents a valuable but underutilized resource in current PRVR approaches. We propose to exploit these underlying cross-modal relationships.
To achieve this, we propose the Inter Correlation Enhancement (ICE) module. Technically, we aim to additionally learn these unlabeled but correlated visual-text relationships, thus constructing a richer semantic space. We focus on exploiting moment-text pairs that are most likely to be semantically corresponding while filtering out those semantically irrelevant ones. The pseudo label assignment process is shown in Figure 3.
Figure 3. The ICE module employs a two-stage selection process for pseudo label assignment. It first computes the similarity between unpaired video moments and text features in a mini-batch. The pairs with mutual maximum similarity are selected as candidate pairs. Then, only those pairs with a similarity higher than the threshold are retained, ensuring high-confidence pseudo labels for subsequent training.
Given a mini-batch with $n$ video-text pairs, we first extract text and moment features as defined in Section 3.2. This results in $n_{m}=nT_{m}$ moment features and $n$ text features. Then, we calculate a similarity matrix ${S}=\left\{s_{ij}\right\}\in\mathbb{R}^{n_{m}\times n}$ for each moment feature and text feature. Note that we manually set the similarities between each text and moments from its paired video to -1 as we aim to mine unpaired correlated moment-text pairs. For the $\bar{i}$-th moment and $\bar{j}$-th text, we find the most similar text and moment, respectively:
(8) $~{}\hat{j}=\mathop{\arg\max}\limits_{j\in\{1,\cdots,n\}}(\{s_{\bar{i}j}\}),~{}\hat{i}=\mathop{\arg\max}\limits_{i\in\{1,\cdots,n_{m}\}}(\{s_{i\bar{j}}\}),$
if $\hat{i}=\bar{i}$ and $\hat{j}=\bar{j}$, they are mutually the most similar to each other and we deem them as a candidate correlated pair.
Intuitionally, a truly correlated pair should not only mutually be the most similar to each other, but also exhibit high similarity. To further filter out mismatched noisy pairs, we pre-define $\beta_{t}$ as the correlation threshold. Only the pairs with a similarity higher than $\beta_{t}$ are selected.
Finally, we obtain $n_{c}$ correlated pairs through the above two-stage selection. These pairs then form a mini-batch of size $n_{c}$ and we align them using contrastive learning losses defined in Equation (5) and (6). The loss of ICE is denoted as $\mathcal{L}_{ice}$.
3.4. Intra Redundancy Mining
The ICE module effectively harnesses inter-sample correlations to enhance the cross-modal semantic space for retrieval. However, while certain video semantics may align with unpaired text queries, they represent redundant information with respect to the originally paired text. This redundancy can introduce noise during video-text matching, thereby compromising retrieval accuracy.
To mitigate the intra-sample redundancy, we propose the Intra Redundancy Mining (IRM) module, which identifies redundant video moments by analyzing relationships among text, video, and moment-level features. These redundant features are then employed as hard negative samples during training. Unlike conventional negative samples (e.g., randomly sampled unpaired videos), these mined negatives exhibit higher feature similarity to the target moment, presenting a more challenging discrimination task. By forcing the model to distinguish these hard negatives, we promote the learning of more discriminative and query-relevant video representations, improving retrieval performance.
We construct redundant video features from two different perspectives. Technically, we model the semantic variation between the target moment feature $m^{t}$, video feature $v$, and query feature $q$ to obtain the video-view redundant feature $r_{v}\in\mathbb{R}^{D}$ and query-view redundant feature $r_{q}\in\mathbb{R}^{D}$:
(9) $r_{v}=FC(v-m^{t}),~{}r_{q}=FC(v-q),$
where $FC$ is a fully-connected layer to bridge the gap between video-moment and video-query differences. The intuition is that the video-level feature $v$ contains global semantics of all moments within the video. By subtracting the target moment feature $m^{t}$ from the video-level feature $v$, we obtain the video-view redundant feature $r_{v}$ which is irrelevant to the text. Similarly, we use a symmetrical operation to obtain the query-view redundant feature $r_{q}$ which is regarded to be irrelevant to the target moment. Since both redundant features $r_{v}$ and $r_{q}$ should be semantically irrelevant to the target moment feature and query feature, we employ $r_{v}$ and $r_{q}$ as negative samples in the moment-level similarity learning (Section 3.2.3), where the loss function is denoted as $\mathcal{L}_{neg}$.
Moreover, we further propose to directly align $r_{v}$ and $r_{q}$. As the two redundant features are obtained at the video-view and query-view respectively, their alignment boosts the learning of underlying relationships between videos and text queries in the cross-modal semantic space. The corresponding loss $\mathcal{L}_{red}$ is also the jointly application of Equation (5) and (6). The overall training objective of this module is the sum of the above two losses:
(10) $\mathcal{L}_{irm}=\mathcal{L}_{neg}+\mathcal{L}_{red}.$
Table 1. Comparison results on TVR, ActivityNet Captions, and Charades-STA. The best results are highlighted in bold, while the second-best outcomes are underlined. We achieve state-of-the-art results on all of the three datasets across all metrics.
Method Venue TVR ActivityNet Captions Charades-STA
R@1 R@5 R@10 R@100 SumR R@1 R@5 R@10 R@100 SumR R@1 R@5 R@10 R@100 SumR
HGR (Chen et al., 2020) CVPR’2020 1.7 4.9 8.3 35.2 50.1 4.0 15.0 24.8 63.2 107.0 1.2 3.8 7.3 33.4 45.7
W2VV++ (Li et al., 2019) MM’2019 5.0 14.7 21.7 61.8 103.2 5.4 18.7 29.7 68.8 122.6 0.9 3.5 6.6 34.3 45.3
DE (Dong et al., 2019) CVPR’2019 7.6 20.1 28.1 67.6 123.4 5.6 18.8 29.4 67.8 121.7 1.5 5.7 9.5 36.9 53.7
DE++ (Dong et al., 2021) TPAMI’2021 8.8 21.9 30.2 67.4 128.3 5.3 18.4 29.2 68.0 121.0 1.7 5.6 9.6 37.1 54.1
RIVRL (Dong et al., 2022b) TCSVT’2022 9.4 23.4 32.2 70.6 135.6 5.2 18.0 28.2 66.4 117.8 1.6 5.6 9.4 37.7 54.3
Cap4Video (Wu et al., 2023) CVPR’2023 10.3 26.4 36.8 74.0 147.5 6.3 20.4 30.9 72.6 130.2 1.9 6.7 11.3 45.0 65.0
XML (Lei et al., 2020) ECCV’2020 10.0 26.5 37.3 81.3 155.1 5.3 19.4 30.6 73.1 128.4 1.6 6.0 10.1 46.9 64.6
ReLoCLNet (Zhang et al., 2021b) SIGIR’2021 10.7 28.1 38.1 80.3 157.1 5.7 18.9 30.0 72.0 126.6 1.2 5.4 10.0 45.6 62.3
CONQUER (Hou et al., 2021) MM’2021 11.0 28.9 39.6 81.3 160.8 6.5 20.4 31.8 74.3 133.1 1.8 6.3 10.3 47.5 66.0
MS-SL (Dong et al., 2022a) MM’2022 13.5 32.1 43.4 83.4 172.4 7.1 22.5 34.7 75.8 140.1 1.8 7.1 11.8 47.7 68.4
JSG (Chen et al., 2023) MM’2023 - - - - - 6.8 22.7 34.8 76.1 140.5 2.4 7.7 12.8 49.8 72.7
UMT-L (Li et al., 2023) ICCV’2023 13.7 32.3 43.7 83.7 173.4 6.9 22.6 35.1 76.2 140.8 1.9 7.4 12.1 48.2 69.6
PEAN (Jiang et al., 2023) ICME’2023 13.5 32.8 44.1 83.9 174.2 7.4 23.0 35.5 75.9 141.8 2.7 8.1 13.5 50.3 74.7
InternVideo2 (Wang et al., 2024a) ECCV’2024 13.8 32.9 44.4 84.2 175.3 7.5 23.4 36.1 76.5 143.5 1.9 7.5 12.3 49.2 70.9
GMMFormer (Wang et al., 2024b) AAAI’2024 13.9 33.3 44.5 84.9 176.6 8.3 24.9 36.7 76.1 146.0 2.1 7.8 12.5 50.6 72.9
BGM-Net (Yin et al., 2024) TOMM’2024 14.1 34.7 45.9 85.2 179.9 7.2 23.8 36.0 76.9 143.9 1.9 7.4 12.2 50.1 71.6
DL-DKD (Dong et al., 2023) ICCV’2023 14.4 34.9 45.8 84.9 179.9 8.0 25.0 37.5 77.1 147.6 - - - - -
Ours - 17.5 39.0 49.9 87.6 194.0 10.1 28.6 41.9 79.8 160.4 2.9 9.2 14.9 54.3 81.3
3.5. Temporal Coherence Prediction
Figure 4. The pipeline of the Temporal Coherence Prediction (TCP) module. The frame features are first divided into distinct groups, and group labels are assigned accordingly. Then, a subset of frame features is randomly selected and shuffled, and the group labels with these frames are predicted.
While our ICE and IRM modules effectively leverage inter-sample correlations and mitigate intra-sample redundancy, their performance fundamentally depends on the model’s ability to discern discriminative video features. Clearer feature distinctions enable more accurate cross-modal correlation measurement in ICE and better redundancy identification in IRM. To further enhance these capabilities, we introduce the Temporal Coherence Prediction (TCP) module, which strengthens temporal feature understanding through self-supervised learning. The pipeline is shown in Figure 4.
Specifically, the TCP module incorporates an auxiliary sequence prediction task. To capture the temporal structure of videos, we leverage the continuity between adjacent frames to partition the video sequence into coherent temporal segments. We group the video features and shuffle them to form an unordered sequence. The model is trained to predict the original temporal order. By explicitly modeling temporal coherence, this task helps the model to learn the coherence and causality between video events. Note that this strategy is applied to both the video-level branch and moment-level branch in Section 3.2.2 to enable multi-grained temporal coherence learning. We mainly take the video-level branch for explanation.
Concretely, we first group frame features into $g$ groups and each frame feature is given its group label ranging from $1$ to $g$. Every $T_{f}/g$ consecutive frame features are assigned with the same group label. Group labels of the original encoded frame features $V_{f}$ are denoted as $y_{o}\in\mathbb{R}^{T_{f}}$. We then randomly select and shuffle some frame features and encode them through the video encoder. Note that shuffling a small or large length of frames might lead to suboptimal performance. Too few shuffled frames provide insufficient learning signals, while excessive shuffling disrupts temporal coherence and makes the task overly challenging. Therefore, we pre-define $\beta_{r}$ as the shuffle ratio. The shuffled frame features and their shuffled group labels are represented as $\hat{V_{f}}=\left\{\hat{f_{i}}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$ and $y_{s}\in\mathbb{R}^{T_{f}}$, respectively. Then we adopt a classifier to predict the group labels of both the original and shuffled frame features $V_{f}$ and $\hat{V_{f}}$:
(11) $p_{o}=Softmax(CLS(V_{f})),~{}p_{s}=Softmax(CLS(\hat{V_{f}})),$
where $CLS$ is a linear classifier, $p_{o}\in\mathbb{R}^{T_{f}\times g}$ and $p_{s}\in\mathbb{R}^{T_{f}\times g}$ are the predicted probability distributions of group labels. We adopt the cross-entropy loss to optimize this module:
(12) $\mathcal{L}_{tcp}=f_{CE}(p_{o},y_{o})+f_{CE}(p_{s},y_{s}),$
where $f_{CE}$ represents the cross-entropy loss function.
3.6. Training and Inference
Our proposed method is optimized by the loss of each module:
(13) $\mathcal{L}=\mathcal{L}_{base}+\lambda_{3}\mathcal{L}_{ice}+\lambda_{4}\mathcal{L}_{irm}+\lambda_{5}\mathcal{L}_{tcp},$
where $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are hyperparameters to balance losses.
During inference, the similarity of the given video-text pair $\mathcal{V}$ and $\mathcal{T}$ is the weighted sum of moment-level and video-level similarities following (Dong et al., 2022a; Wang et al., 2024b; Dong et al., 2023; Jiang et al., 2023; Yin et al., 2024):
(14) $\mathcal{S}\left(\mathcal{V},\mathcal{T}\right)=\alpha~{}\mathcal{S}_{m}\left(\mathcal{V},\mathcal{T}\right)+\left(1-\alpha\right)\mathcal{S}_{v}\left(\mathcal{V},\mathcal{T}\right),$
where $\alpha\in\left[0,1\right]$ is a hyperparameter to balance the similarities.
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets
To evaluate the effectiveness of our method on the PRVR task, we utilize three untrimmed video datasets: TVR (Lei et al., 2020), ActivityNet Captions (Krishna et al., 2017), and Charades-STA (Gao et al., 2017). Note that we do not use moment-level annotations provided by these datasets in the setting of PRVR. Unlike conventional T2VR datasets (Xu et al., 2016; Chen and Dolan, 2011; Wang et al., 2019) using pre-trimmed short clips, PRVR employs untrimmed long videos where the content is only partially relevant to text queries.
TVR is collected from six TV shows, containing 21.8K videos and 109K text queries. Each video has five textual descriptions on average, depicting different moments in the video. Following (Dong et al., 2022a), the training set consists of 17,435 videos with 87,175 moments, while the test set contains 2,179 videos with 10,895 moments.
ActivityNet Captions consists of a wide range of human activities, sourced from 20K YouTube videos. On average, each video has 3.7 moments with sentence descriptions. We adopt the same data partition as in (Dong et al., 2022a, 2023; Wang et al., 2024b), where 10,009 and 4,917 videos (37,421 and 17,505 moments) are utilized as the training and test sets.
Charades-STA primarily comprises videos that depict indoor activities, including 6,670 videos and 16,128 moment annotations. We utilize the official data partition, 12,408 and 3,720 moment-sentence pairs for training and evaluation, respectively.
4.1.2. Evaluation Metrics
Following previous works (Dong et al., 2022a, 2023; Wang et al., 2024b), we adopt the rank-based metrics Recall Rate at $k$ ($R@k$) and the sum of all recall rates ($SumR$). $R@k$ measures the proportion of relevant items successfully retrieved within the top-$k$ results, with $k$ set to {1, 5, 10, 100}. Higher recall rates indicate better retrieval accuracy.
4.1.3. Implementation Details
To ensure a fair comparison with previous works on PRVR (Dong et al., 2022a; Jiang et al., 2023; Dong et al., 2023; Wang et al., 2024b; Yin et al., 2024), we utilize the same visual and textual features provided by (Dong et al., 2022a). The visual and textual features are exacted by pre-trained I3D (Carreira and Zisserman, 2017) and RoBERTa (Liu et al., 2019), respectively. The feature dimension $D$ is set to 384. $T_{m}$, $m$, and $\alpha$ are set to 32, 0.2, and 0.7 following (Dong et al., 2022a; Wang et al., 2024b). $\beta_{t}$ and $\beta_{r}$ are pre-defined as 0.4 and 0.25. $g$ is set to 8. We set $\lambda_{\{1,2,3,4,5\}}$ to {0.02, 0.04, 0.1, 1, 1}. For model training, we utilize the Adam optimizer (Kingma and Ba, 2015) with an initial learning rate of 2.5e-4 and a mini-batch size of 128. The model is trained for 100 epochs with an early stopping strategy. Dropout (Srivastava et al., 2014) is set to 0.15 to prevent overfitting. All experiments are conducted on a single NVIDIA RTX 3090 GPU.
4.2. Comparison with State-of-the-Arts
We compare the proposed method with other state-of-the-art methods on TVR, ActivityNet Captions, and Charades-STA. These methods consist of both PRVR and T2VR models. The results are borrowed from (Dong et al., 2022a; Wang et al., 2024b) and the reported results of corresponding papers.
As summarized in Table 1, our method achieves state-of-the-art (SOTA) performance on all three datasets across all metrics. Specifically, we achieve significant improvements in SumR across these benchmarks: (1) On the TVR dataset, we outperform the previous SOTA method DL-DKD (Dong et al., 2023) by 14.1% absolute and 7.84% relative improvements. (2) On ActivityNet Captions, we surpass DL-DKD with gains of 12.8% absolute and 8.67% relative improvements. (3) On Charades-STA, we exceed PEAN (Jiang et al., 2023) by 6.6% absolute and 8.84% relative improvements. These results demonstrate the effectiveness and robustness of our approach.
4.3. Ablation Study
4.3.1. Main Ablation Studies
Table 2. Main ablation studies of the proposed key modules on Charades-STA. The best results are in bold.
Row Setting R@1 R@5 R@10 R@100 SumR
ICE IRM TCP
1 1.5 6.8 11.3 49.5 69.1
2 2.2 7.5 12.4 52.2 74.3
3 2.1 7.8 12.9 51.9 74.7
4 1.7 7.0 12.0 50.5 71.2
5 2.4 8.3 13.8 52.1 76.6
6 2.5 8.5 13.6 52.2 76.8
7 2.6 8.8 14.2 52.6 78.2
8 2.9 9.2 14.9 54.3 81.3
Table 3. Ablation study of loss terms in the Intra Redundancy Mining module on Charades-STA.
Row Loss Terms R@1 R@5 R@10 R@100 SumR
$\mathcal{L}_{neg}$ $\mathcal{L}_{red}$
1 2.5 8.5 13.6 52.2 76.8
2 2.3 8.9 14.9 53.2 79.3
3 2.3 8.7 14.7 52.8 78.5
4 2.9 9.2 14.9 54.3 81.3
To comprehensively validate the effectiveness of each module in our method, we conduct extensive ablation studies in Table 2. In particular, Row 1 indicates the base model trained with $\mathcal{L}_{base}$ only, whose retrieval accuracy is far below the full model (Row 8) trained with Equation (13). Comparing the results in each row, we have the following observations:
(1) Each module independently improves performance over the base model (Row 1). The ICE module (Row 2) boosts SumR by 5.2, demonstrating its ability to enhance inter-sample semantic alignment. The IRM module (Row 3) improves SumR by 5.6, indicating its role in refining the cross-modal semantic space. The TCP module (Row 4) achieves the gain of 2.1 in SumR, highlighting its effectiveness in learning discriminative visual features.
(2) Combining any two modules yields further improvements (Row 5-7), indicating their complementary roles in aligning semantics. Specifically, the joint application of ICE and IRM achieves the largest gain (Row 7), highlighting the necessity of exploiting the cross-modal dual nature of video-text modalities in PRVR.
(3) Integrating all three modules leads to the best performance (Row 8), with significant gains across all metrics. This proves the effectiveness of jointly leveraging inter-sample correlations, intra-sample redundancy mining, and temporal coherence learning.
4.3.2. Inter Correlation Enhancement
We validate the efficacy of the correlation threshold $\beta_{t}$ in the Inter Correlation Enhancement module. As presented in Figure 5 (a), the choice of $\beta_{t}$ affects retrieval performance. When $\beta_{t}$ is set to 0.4, the model achieves the highest SumR of 81.3, outperforming both lower and higher thresholds. Lowering $\beta_{t}$ to 0.2 increases noise from weakly correlated pairs, reducing SumR to 80.1. On the other hand, raising $\beta_{t}$ to 0.6 or 0.8 over-suppresses potential correlations, thus degrading SumR to 79.6 and 78.6, respectively. Notably, the absence of $\beta_{t}$ (i.e., retaining all candidate pairs) yields a suboptimal SumR of 79.1, demonstrating that indiscriminate inclusion of low-similarity pairs could introduce harmful noise during training.
Figure 5. Ablation study of the correlation threshold $\beta_{t}$ in the Inter Correlation Enhancement module and the shuffle ratio $\beta_{r}$ in the Temporal Coherence Prediction module.
4.3.3. Intra Redundancy Mining
We investigate the effectiveness of each loss function and their combinations in the Intra Redundancy Mining (IRM) module. As shown in Table 3, both $\mathcal{L}_{neg}$ (Row 2) and $\mathcal{L}_{red}$ (Row 3) lead to improvements across all metrics. Moreover, combining $\mathcal{L}_{neg}$ and $\mathcal{L}_{red}$ yields the highest SumR, outperforming either loss alone. The effectiveness arises because $\mathcal{L}_{neg}$ directly suppresses redundancy by pushing away hard negatives, while $\mathcal{L}_{red}$ implicitly strengthens cross-modal alignment by correlating video-view and query-view redundancies.
4.3.4. Temporal Coherence Prediction
We assess the impact of the shuffle ratio $\beta_{r}$ in the Temporal Coherence Prediction (TCP) module. As illustrated in Figure 5 (b), the choice of $\beta_{r}$ influences the model’s ability to learn temporal coherence. When $\beta_{r}=0.25$, the model achieves the optimal SumR of 81.3, outperforming all other configurations. Adopting other ratios results in degradations in performance: smaller ratios yield insufficient learning signals, as minimal shuffling fails to challenge the model’s temporal reasoning. Conversely, larger ratios excessively disrupt temporal coherence, making the reordering task overly ambiguous and harming feature learning. Notably, the worst $\beta_{r}$ configuration even underperforms the full absence of TCP, demonstrating the necessity of setting a proper shuffle ratio.
We then demonstrate the contribution of applying the TCP module to both the video-level branch and the moment-level branch. As depicted in Table 4, leveraging TCP solely on the video-level branch (Row 2) improves SumR to 79.7, primarily enhancing fine-grained temporal modeling. Besides, applying TCP only to the moment-level branch (Row 3) achieves an improved SumR of 80.2, as moment-level coherence learning encourages the encoder to reason coarse-grained temporal relationships. Furthermore, the full setup (Row 4) integrates these advantages, achieving the best performance across all metrics. This dual-branch design ensures holistic temporal coherence learning at multiple granularities, which is essential for robust temporal feature learning.
4.3.5. Base Model
To demonstrate the broader generalization of our method, we further integrate it into two existing PRVR models: MS-SL (Dong et al., 2022a) and GMMFormer (Wang et al., 2024b). Concretely, we directly replace the base model with these models. As shown in Table 5, our approach acts as a plug-and-play enhancement to these models on the Charades-STA dataset. Consistent performance improvements are observed across both models. For instance, our method boosts SumR by 16.96% for MS-SL and 13.72% for GMMFormer in relative improvements, demonstrating its effectiveness and generalization. These results suggest that existing PRVR models inherently suffer from the limited cross-model alignment. In contrast, our method systematically exploits the inter-sample correlation and intra-sample redundancy, resulting in substantial performance improvements. The universal gains also highlight that our method is model-agnostic and effectively constructs a more comprehensive cross-modal semantic space.
Table 4. Ablation study of the Temporal Coherence Prediction module on Charades-STA.
Row Setting R@1 R@5 R@10 R@100 SumR
Video Moment
1 2.6 8.8 14.2 52.6 78.2
2 2.7 9.0 14.6 53.4 79.7
3 2.8 8.9 14.8 53.7 80.2
4 2.9 9.2 14.9 54.3 81.3
Table 5. Effect on different base models on Charades-STA. We directly apply our proposed three modules to existing PRVR methods.
Method Setting R@1 R@5 R@10 R@100 SumR
MS-SL (Dong et al., 2022a) Origin 1.8 7.1 11.8 47.7 68.4
+ Ours 2.4 9.1 14.6 53.9 80.0
GMMFormer (Wang et al., 2024b) Origin 2.1 7.8 12.5 50.6 72.9
+ Ours 3.0 9.5 15.3 55.1 82.9
Figure 6. Visualization comparisons of retrieval results between our method, MS-SL (Dong et al., 2022a), and GMMFormer (Wang et al., 2024b) on ActivityNet Captions. Top-5 retrieval results are shown from left to right. Ground-truth videos are marked with red boxes. Zoom in for better visibility.
(a) Base Model
(b) Full Model (Ours)
Figure 7. t-SNE visualization (van der Maaten and Hinton, 2008) of text and video features on Charades-STA. (a) is the base model trained with only $\mathcal{L}_{base}$. (b) shows the full model trained with Equation (13).
4.4. Qualitative Results
4.4.1. Retrieval Results
To qualitatively validate the effectiveness of our method, we report visualization comparisons of retrieval results between our method, MS-SL (Dong et al., 2022a), and GMMFormer (Wang et al., 2024b) on ActivityNet Captions in Figure 6. In the first case, MS-SL and GMMFormer superficially match videos with the action riding and mistakenly associate riding with bikes, ignoring key semantics including pushes a rake and scooter. In contrast, our method accurately retrieves the target video with correct semantics. The remaining example also shows a similar phenomenon, which demonstrates the robust and superior retrieval capability of our method.
4.4.2. t-SNE Visualization
We report t-SNE visualization (van der Maaten and Hinton, 2008) to analyze the alignment of video and text features in the semantic space. For clearer observation, we randomly sample a subset of paired videos and text queries from Charades-STA. As shown in Figure 7 (a), in the base model, the distributions of video and text features exhibit limited overlap in the shared cross-modal space, indicating weak alignment between the two modalities. In contrast, our method achieves a more discriminative feature distribution. As in Figure 7 (b), video and text features are tightly interleaved in regions and the regions maintain sufficient separation to preserve semantics-specific distinctions. This suggests that our approach successfully aligns the two modalities while retaining their unique characteristics, leading to improved semantic alignment for PRVR.
5. Conclusion
In this paper, we present a novel framework for Partially Relevant Video Retrieval, which systematically exploits the cross-model dual nature of video-text modalities: inter-sample correlation and intra-sample redundancy. Our approach introduces three key modules. The Inter Correlation Enhancement module strengthens cross-modal alignment by identifying and leveraging semantically similar yet unpaired text-video samples as pseudo-positive pairs. The Intra Redundancy Mining module improves discriminative learning by treating redundant video moments as hard negative samples, forcing the model to focus on query-relevant visual semantics. The Temporal Coherence Prediction module enhances temporal feature discrimination through a self-supervised video sequence prediction task. Extensive experiments on three datasets demonstrate that our framework achieves state-of-the-art performance.